Problem 5-22 CVP Applications; Contribution Margin Ratio; Break-Even Analysis; Cost Structure (LO5-1, LO5-3, LO5-4, LO5-5, LOProblem 5-22 CVP Applications; Contribution Margin Ratio; Break-Even Analysis; Cost Structure (LO5-1, LO5-3, LO5-4, LO5-5, LOProblem 5-22 CVP Applications; Contribution Margin Ratio; Break-Even Analysis; Cost Structure (L05-1, LO5-3, LO5-4, LO5-5, LOProblem 5-22 CVP Applications; Contribution Margin Ratio; Break-Even Analysis; Cost Structure (LO5-1, LO5-3, LO5-4, LO5-5, LOProblem 5-22 CVP Applications; Contribution Margin Ratio; Break-Even Analysis; Cost Structure (LO5-1, LO5-3, LO5-4, LO5-5, LOProblem 5-22 CVP Applications; Contribution Margin Ratio; Break-Even Analysis; Cost Structure (LO5-1, LO5-3, LO5-4, LO5-5, LOProblem 5-22 CVP Applications; Contribution Margin Ratio; Break-Even Analysis; Cost Structure (LO5-1, LO5-3, LO5-4, LO5-5, LO
Answer #1

Solution 1:

CM ratio = Contribution margin / Sales = $160,800 / $402,000 = 40%

break-even point in unit sales = Fixed expenses / CM per unit = $178,800 / $8 = 22350 units

Breakeven point in dollar sales = Fixed expenses / CM ratio = $178,800 / 40% = $447,000

Solution 2:

increase (decrease) in the company’s monthly net operating income = Increase in contribution margin - Increase in advertising expenses

= ($80,000*40%) - $6,500 = $25,500

Solution 3:

New selling price per unit = $30*90% = $27

New CM per unit = $12 - $3 = $9 per unit

New fixed costs =$178,800 + $33,000 = $211,800

New sales volume = 13400*2 = 26800 units

New operating income = Contribution margin - Fixed costs = (26800*$9) - $211,800 = $29,400

Solution 4:

New contribution margin per unit = $12 - $0.60 = $11.40 per unit

Target income = $4,600

Nos of units to be sold to attain target profit = (Fixed cost + Net operating income) / CM per unit

= ($178,800 + $4,600) / $11.40 = 16088 units

Solution 5a:

New CM per unit = $12 +$3 = $15 per unit

New fixed costs = $178,800 + $53,000 = $231,800

New CM ration = $15 / $30 = 50%

New break even point in units = $231,800 / $15 = 15453 units

Breakeven point in dollar sales = $231,800 / 50% = $463,600

Solution 5b:

Contribution format income statement - Operation Automated
Particulars Per unit % Total
Sales $30.00 100% $600,000.00
Variable costs $15.00 50% $300,000.00
Contribution margin $15.00 50% $300,000.00
Fixed costs $231,800.00
Net operating income $68,200.00
Contribution format income statement - Operation not Automated
Particulars Per unit % Total
Sales $30.00 100% $600,000.00
Variable costs $18.00 60% $360,000.00
Contribution margin $12.00 40% $240,000.00
Fixed costs $178,800.00
Net operating income $61,200.00

Solution 5c:

As net income is increasing, therefore company should automate its operations.